# A Novel Federated Learning-Based IDS for Enhancing UAVs Privacy and Security

### v2

@@ VI-B Communication and Response
While C-IDS demonstrates superior detection performance, it achieves this by collecting local data from each node periodically in order to have a full picture of the network. However, this comes with a communication cost. For C-IDS, the communication cost (CC) for each period can be calculated using the formula:
$CC=N*F*S$
where N is the total number of clients, F is the number of features collected and sent from each client in each time interval, S is the average size of each feature. In our experiments, with 50 clients/nodes in the network and each node collects 31 features, with each feature comprising 4 bytes. The equation calculates the communication cost in one period of data collection. In our simulation, considering that data is collected every 5 s over a duration of 1800 s, the total cost is computed by multiplying ${1800}/{5}=360$. In tasks where the duration extends or UAVs collect data at the start of each task, the cost will increase proportionally with the duration. Moreover, as the network faces new types of attacks, the number of collected features might need to be expanded. For example, to detect dropping attacks and distinguish malicious dropping from legitimate dropping due to the use of wireless links, lower-layer features might need to be included [62].
In contrast, FL-IDS only collects weights from each client to contribute the global model during (re-)training. Although it requires every weight value in the DNN arhcitecture to be sent periodically by all nodes or the best client nodes (in the BTSC approach), such information is only needed during training. Moreover, since FANETs might experience high packet loss due to high mobility, FL’s capacity to handle missing data and adapt provides the development of robust models [52].
Furthermore, in comparison to C-IDS, FL-IDS offers the advantage of quicker response times to attacks. This is attributed to the decentralized nature of FL, which eliminates the need for periodic data transmission to a central server. As a result, FL-based IDS can promptly detect and respond to threats without the latency associated with continuous data transmission and centralized processing.
VI-C Resource Consumption
In the proposed model, energy is an important issue to consider, especially during the training process. The UAVs train their models locally using onboard datasets, ensuring data privacy without transmitting raw data to a central server. This decentralized training approach, while advantageous for privacy, does incur energy costs on the UAVs. The energy consumption during training is influenced by factors such as the model’s complexity, dataset size, federated learning implementation strategies, and the computational capabilities of the UAV. To mitigate this, data collection intervals can be reduced, and more energy-efficient machine learning techniques can be incorporated. While ensuring accurate model updates remains a priority, solutions should be mindful of the potential trade-offs between achieving accuracy and managing energy consumption.
The BTSC strategy can be feasible, as shown in the results. Selecting the best detection nodes and only their participation in the training process will save the energy of non-participant UAVs. Split learning technique with FL [63] can be another promising solution to address energy consumption limitations. This approach involves a practical solution in this scenario, where only a small part of the model is kept and trained on clients, while the larger portion of the model remains on the servers. Furthermore, few-shot learning with FL proposes a new avenue for energy saving with the use of few data for training [64].
While each node in L-IDS consumes energy similar to each node in FL-IDS, the lack of communication between nodes in L-IDS necessitates that each node maintains its own model. Conversely, with the use of the BTSC strategy in FL-IDS, the model can be distributed among only select nodes. However, C-IDS incurs higher energy consumption due to the periodic data transmission and processing of large amounts of data at the GBS. This transmission, especially in wireless nodes and UAV devices, significantly contributes to overall energy expenditure [65]. Hence communication also plays a substantial role in energy consumption. Furthermore, storage costs might pose an issue for C-IDS, given its periodic collection of large data from each node. Last but not the least, environmental factors such as wind and other physical elements may impact energy usage of UAVs. Wind direction and speed are factors that affect UAV energy consumption. Flying with the wind can reduce energy usage, while facing headwinds can increase power consumption due to the additional thrust required [66]. In addition to wind effects, extreme cold and hot temperatures also impact the operation of UAVs. For example, cold temperatures can adversely affect UAV batteries [67].
VI-D Privacy and Security
@@
While Central IDS (C-IDS) generally exhibits high performance in detecting attacks, it necessitates the central collection of local data, which raises significant concerns about data privacy and security. This collected data could contain sensitive information, such as communication content, operation-specific or node-specific information, and location data. Additionally, a single C-IDS system can become an attractive target for potential attackers. If the system is compromised, attackers could gain access to all this sensitive data.
@@
Even though FL has a lot of potential benefits and provide privacy, it still could be susceptible to adversarial attacks, much like other types of ML methods. Adversaries may attempt to undermine the integrity of the FL process by providing erroneous or misleading data. This could involve intentionally supplying data that is inaccurate or deceptive, with the goal of misleading the trained model. Additionally, adversaries might engage in the malicious engineering of data, where they intentionally manipulate data in a deceptive manner to trick the trained model into making mistakes [68].
@@
While adversarial attacks is a common issue for all types of IDS; C-IDS and F-IDS also introduce a single point of failure due to collecting and/or processing data at the GBS. On the other hand, while the global model can be moved to another node, C-IDS requires the existence of a powerful node due to processing and storing large amount of data. In the case of L-IDS, where each node operates independently, the decentralized approach limits the potential impact of a compromise to individual nodes. However, the lack of inter-node communication may also restrict the system’s ability to implement collective defense mechanisms or coordinated responses to sophisticated attacks, potentially limiting its overall resilience against certain types of threats.
VII Conclusion
@@
This study introduces a novel approach to intrusion detection in FANETs by proposing Federated IDS (FL-IDS). Given the distributed nature of network data in FANETs, a distributed and cooperative architecture is inherently more suitable. However, in such an architecture, data sharing raises privacy concerns, particularly in the context of UAVs engaged in mission-critical applications. FL-IDS offers a solution to address these concerns, and in this study, it is introduced and comprehensively compared with traditional approaches: C-IDS and L-IDS.
@@
The comparative evaluation covers multiple metrics, including effectiveness, communication and response, resource consumption, privacy and security considerations. The assessment extends to various networks with varying mobility patterns to ensure a comprehensive analysis. Moreover, our study incorporates realistic network scenarios that account for factors such as 3D node movement, local data collection by each node, and realistic traffic patterns. To the best of our knowledge, this study represents one of the few recent studies in the field of FANET security [18][20] where data is based on an actual FANET dataset, but with more realistic settings specifically tailored to suit the dynamics and requirements of FANETs. In previous studies [41], the application of federated learning was often demonstrated using synthetic data that mimicked IoT data federations. We believe this study contributes significantly to the field as it represents the first comprehensive exploration of the use of federated learning for routing attacks and offers a detailed comparison with traditional intrusion detection approaches. The experimental results demonstrate the effectiveness and suitability of the proposed FL-IDS approach driven by its distributed training methodology, while also addressing critical privacy concerns.

### v3

@@ 6.2 Communication and Response
While C-IDS demonstrates superior detection performance, it achieves this by periodically collecting local data from each node to have a comprehensive view of the network. However, this comes with a communication cost. For C-IDS, the communication cost (CC) for the training process can be calculated using the formula:
$CC=N*F*S*Periods$
where N is the number of UAVs (clients), F is the number of features collected per UAV, S is the size of each feature in bytes, and Periods is the number of data collection intervals. In our experiments, the network consisted of 50 clients/nodes, each collecting 31 features, with each feature occupying 4 bytes. Data were collected over a total duration of 1800 seconds, consisting of 360 collection periods, each lasting 5 seconds, during the benign-only scenario. However, when attackers are included, the total number of collection periods doubles to 720, resulting in a total data collection duration of 3600 seconds. The communication cost is calculated as CC = 4,464,000 bytes. This cost will increase proportionally with the duration of the operation or if additional features are required to address new types of attacks, such as lower layer features to detect malicious packet drop [64].
In contrast, FL-IDS only collects weights from each client to contribute the global model during (re-)training. Although it requires every weight value in the model architecture periodically, the BTSC strategy (e.g., the top 20% based on detection accuracy) further minimizes this cost by limiting training participation to a subset of the most effective nodes, thus reducing communication overhead. The communication cost can be expressed as:
$CC_{F}=2*N*W*S*Epoch$
where N is the number of participating UAVs, W represents the weights of the transmitted model, S is the size of each weight in bytes and Epoch is the number of the training cycle. The factor of 2 accounts for both uploading the local model updates to the server and downloading the global model from the server. In our experiment, $CC_{F}$ is calculated as 3,336,000 byte for BTSC.
FL-IDS can result in higher costs when using complex models, primarily due to the transmission of large model weights or the requirement of a greater number of training epochs. However, FL-IDS eliminates the need for continuous data transmission in testing, as required by C-IDS, thereby enhancing efficiency for real-time detection scenarios. FL-IDS also offers inherent advantages for dynamic environments like FANETs. Given the high mobility and packet loss in such networks, FL’s ability to handle missing data during training ensures robust model updates without requiring constant retransmissions [54]. Furthermore, FL-IDS enables faster response times to security threats, as its decentralized structure eliminates the need for frequent data transmission to a central server. This allows real-time threat detection and mitigation without the latency associated with centralized data processing.
6.3 Resource Consumption
Energy efficiency is a critical consideration in FL-IDS, particularly during the training process. UAVs perform local training using onboard datasets, ensuring data privacy by avoiding raw data transmission to a central server. While this approach enhances privacy, it can cause computational and energy costs, which depend on factors such as model complexity, dataset size, and computational capabilities of the UAVs.
To evaluate the energy overhead of the proposed scheme, we qualitatively estimate the per-node energy consumption as the sum of the energy required for local training and the communication energy for sharing model weights during the aggregation process. The energy consumption of communication is directly proportional to the communication cost, which can be represented as:
$E_{com}=W*S*Epoch$
The total energy consumption for each node, denoted as $E_{total}$, is calculated as the sum of the energy consumed during local training ($E_{t}$) and the communication energy ($E_{com}$):
$E_{total}=E_{t}+E_{com}$
To optimize energy consumption in the proposed scheme, strategies can be applied, such as increasing the interval between communication rounds, reducing the frequency of data collection, or employing energy-efficient machine learning techniques. These strategies aim to reduce energy usage while maintaining a balance between conserving energy and achieving high model accuracy. Careful tuning of these parameters is necessary to appropriately manage the trade-off between energy efficiency and model performance.
The BTSC strategy further reduces energy consumption by involving only the best-performing UAVs in training, conserving energy for non-participating nodes. Other promising techniques include split learning [65], where only part of the model is trained locally while the rest resides on a central server, and few-shot learning [66], which minimizes training data requirements.
For comparison, L-IDS requires each node to independently train and maintain its model, with energy consumption represented by $E_{t}$. Unlike FL-IDS, L-IDS does not rely on data transmission to a central server, leading to a reduction in energy overhead. However, FL-IDS with the BTSC strategy distributes the workload more efficiently, reducing energy consumption across the network. In contrast, C-IDS incurs significantly higher energy costs because of the periodic transmission of large data and the centralized processing requirements at the GBS. In our experiment, the energy overhead of C-IDS is calculated based on $CC*\epsilon$. This centralized approach, particularly in wireless UAV networks, contributes substantially to energy expenditure [67]. Furthermore, C-IDS faces storage challenges due to the need to store large volumes of data collected periodically.
6.4 Privacy and Security
@@
While Central IDS (C-IDS) generally exhibits high performance in detecting attacks, it necessitates the central collection of local data, which raises significant concerns about data privacy and security. This data collected could contain sensitive information, such as communication content, operation-specific or node-specific information, and location data. In addition, a single C-IDS system can become an attractive target for potential attackers. If the system is compromised, attackers could gain access to all this sensitive data.
@@
Although FL has a lot of potential benefits and provides privacy, it could still be susceptible to adversarial attacks, much like other types of ML methods. Adversaries may attempt to undermine the integrity of the FL process by providing erroneous or misleading data. This could involve intentionally providing inaccurate or deceptive data with the goal of misleading the trained model. Additionally, adversaries might engage in the malicious engineering of data, where they intentionally manipulate data in a deceptive manner to trick the trained model into making mistakes [68].
To ensure secure model transmission over in-band communication in federated learning-based IDS, several security mechanisms can be employed. First, end-to-end encryption can be implemented to maintain the confidentiality and integrity of models during transmission, preventing adversaries from intercepting or tampering with them [69]. Second, the adoption of blockchain technology can establish a tamper-proof record of model updates, enhancing transparency and allowing nodes to track the history of model changes [70]. Third, robust anomaly detection mechanisms can be deployed to validate received model updates, identifying and discarding malicious or suspicious contributions before they affect the global model [71]. These mechanisms analyze update distributions to detect deviations indicative of poisoning or model manipulation, thus ensuring the integrity of the global model. Additionally, techniques such as secure multiparty computation (SMC) and differential privacy (DP) can further enhance security. SMC allows multiple parties to securely aggregate updates without revealing individual inputs, preserving confidentiality throughout the aggregation process. DP, on the other hand, improves privacy by introducing noise to updates, preventing the identification of individual contributions, and mitigating potential privacy breaches while maintaining model performance.
@@
Although adversarial attacks are a common issue for all types of IDS; C-IDS and F-IDS also introduce a single point of failure due to collecting and/or processing data at the GBS. On the other hand, while the global model can be moved to another node, C-IDS requires the existence of a powerful node due to processing and storing large amount of data. In the case of L-IDS, where each node operates independently, the decentralized approach limits the potential impact of a compromise to individual nodes. However, the lack of inter-node communication may also restrict the system’s ability to implement collective defense mechanisms or coordinated responses to sophisticated attacks, potentially limiting its overall resilience against certain types of threat.
7 Conclusion
@@
This study introduces a novel approach to intrusion detection in FANETs by proposing Federated IDS (FL-IDS). Given the distributed nature of network data in FANETs, a distributed and cooperative architecture is inherently more suitable. However, in such an architecture, data sharing raises privacy concerns, particularly in the context of UAVs engaged in mission-critical applications. FL-IDS offers a solution to address these concerns and, in this study, it is introduced and compared comprehensively with traditional approaches. C-IDS and L-IDS.
@@
The comparative evaluation covers multiple metrics, including effectiveness, communication and response, resource consumption, privacy and security considerations. The assessment extends to various networks with varying mobility patterns to ensure a comprehensive analysis. In addition, our study incorporates realistic network scenarios that account for factors such as 3D node movement, local data collection by each node, and realistic traffic patterns. To our knowledge, this study represents one of the few recent studies in the field of FANET security [18][20] where data are based on an actual FANET dataset, but with more realistic settings specifically tailored to suit the dynamics and requirements of FANETs. In previous studies [43], the application of federated learning was often demonstrated using synthetic data that mimicked federations of IoT data. We believe that this study contributes significantly to the field as it represents the first comprehensive exploration of the use of federated learning for routing attacks and offers a detailed comparison with traditional intrusion detection approaches. The experimental results demonstrate the effectiveness and suitability of the proposed FL-IDS approach driven by its distributed training methodology, while also addressing critical privacy concerns.